OpenStudy (anonymous):

I asked this kind of question last week and didn't receive the answer I was hoping for.

OpenStudy (anonymous):

The question I have right now is "State whether the given measurements determine zero, one, or two triangles. B = 84°, b = 28, c = 25 "

OpenStudy (anonymous):

How do I know when The 3 measurements can give 1, 2 or no triangles?

OpenStudy (anonymous):

I know a triangle cannot be made when something like a + c < b

jimthompson5910 (jim_thompson5910):

have a look at this article http://www.regentsprep.org/regents/math/algtrig/att12/lawofsinesAmbiguous.htm

OpenStudy (anonymous):

In this article they say: Angles could be 30, 53, and 97 : sum 180 The angle from Quadrant II could create angles 30, 127, and 23 : sum 180 because 180 - 53 = 27?

OpenStudy (anonymous):

I know sin( x - 180) = sin x

OpenStudy (anonymous):

Or is it sin( 180 - x ) ?

jimthompson5910 (jim_thompson5910):

it's sin(180 - x) = sin(x)

OpenStudy (anonymous):

Ok. So for my problem I got A = 84 B = 62.6 C = 33.4

OpenStudy (anonymous):

I dont know where to go from here

jimthompson5910 (jim_thompson5910):

you should have something like sin(C)/c = sin(B)/b sin(C)/25 = sin(84)/28 as part of your steps

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

I'm getting this sin(C)/c = sin(B)/b sin(C)/25 = sin(84)/28 C = arcsin(25*sin(84)/28) C = 62.618755 so I'm not sure how you got C = 33.4

jimthompson5910 (jim_thompson5910):

Drawing 

OpenStudy (anonymous):

I got that too but 84 + 62.6 + x = 180

OpenStudy (anonymous):

I mean A is 33.4

jimthompson5910 (jim_thompson5910):

oh ok, so that's the angle for A (not C) which means A = 33.4 (approximate) B = 84 C = 62.6 (approximate)

jimthompson5910 (jim_thompson5910):

yeah you beat me to it

OpenStudy (anonymous):

How do I know how many triangles can be formed though?

jimthompson5910 (jim_thompson5910):

there's another possible angle for B, but the question is: does it make a triangle? you'll have to consider angle A and C

OpenStudy (anonymous):

Why B?

jimthompson5910 (jim_thompson5910):

oh my bad, I meant C

jimthompson5910 (jim_thompson5910):

there's another possible angle for C, but the question is: does it make a triangle? you'll have to consider angle A

OpenStudy (anonymous):

Why are you picking A?

jimthompson5910 (jim_thompson5910):

ok ignore A for a second

jimthompson5910 (jim_thompson5910):

we found that C = 62.6 (roughly) what is another possible value of C?

OpenStudy (anonymous):

117.4

jimthompson5910 (jim_thompson5910):

so if C = 117.4 (roughly), then what is angle A?

OpenStudy (anonymous):

Im not sure

jimthompson5910 (jim_thompson5910):

we know B = 84

jimthompson5910 (jim_thompson5910):

B = 84 doesn't change

jimthompson5910 (jim_thompson5910):

only A and C are allowed to change

OpenStudy (anonymous):

62.6 - 54.8 = 7.8

jimthompson5910 (jim_thompson5910):

A+B+C = 180 A = unknown B = 84 C = 117.4

OpenStudy (anonymous):

A + B + C doesn't equal 180 then

jimthompson5910 (jim_thompson5910):

if B = 84 and C = 117.4, then what is A? use A+B+C = 180

OpenStudy (anonymous):

117.4 + 84 > 180

OpenStudy (anonymous):

A must be -21.4

jimthompson5910 (jim_thompson5910):

since you can't have a negative angle, that means the other triangle is not possible

OpenStudy (anonymous):

Oh ok

jimthompson5910 (jim_thompson5910):

so only one triangle is possible and that's with A = 33.4 B = 84 C = 62.6

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

np